OpenStudy (ashleyy_smith):

52 decreased by twice Vanessa's age Use the variable v to represent Vanessa's age.

OpenStudy (ashleyy_smith):

@Ghastlyspec

OpenStudy (anonymous):

Of, so you need this made into an equation?

OpenStudy (ashleyy_smith):

yes :o

OpenStudy (anonymous):

Okay, so v=Vanessa's age. We are subtracting 52 by vanessa's age. x=52 - 2v should be your equation.

OpenStudy (anonymous):

Do you understand the equation?

OpenStudy (ashleyy_smith):

x=52? O.o

OpenStudy (ashleyy_smith):

where do I get x from

OpenStudy (anonymous):

Okay so x is the sum, which means 52 minus twice vanessa's age.

OpenStudy (anonymous):

Like 5 + 5=x, x is the unknown variable which in this case is your answer.

OpenStudy (anonymous):

Any other part of it you do not understand?

OpenStudy (anonymous):

@Ashleyy_Smith Any other questions?

OpenStudy (ashleyy_smith):

Yes ((: Aldo is riding his bicycle. He rides at a speed of 9 kilometers per hour for 22.5 kilometers. For how many hours does he ride?

OpenStudy (anonymous):

Okay, every hour is 9 km, We have guaranteed 2 hours since 9*2=18, Then we have about 4.5 km extra.

OpenStudy (anonymous):

We can instantly say there is 1/3 of an hour extra but what about the other 1.5 km?

OpenStudy (anonymous):

We currently have 2 hours and 20 minutes with 21 km traveled.

OpenStudy (ashleyy_smith):

mhm I see

OpenStudy (anonymous):

Now for the last 1.5, If 3 km is 20 minutes, How long is 1.5?

OpenStudy (ashleyy_smith):

ummmm 13?

OpenStudy (anonymous):

Okay, 3 km is 20 minutes, 1.5 + 1= 2.5. Therefore 1.5 + 1.5 = 3

OpenStudy (anonymous):

3k = 20 minutes Divide 3 in half

OpenStudy (ashleyy_smith):

20/3? right?

OpenStudy (anonymous):

No, that is 20 divided by 3. I am asking to divide 3 in half, If you have three pieces of a brownie left and you and one friend are sharing, how would you split it evenly so you both have the same amount?

OpenStudy (ashleyy_smith):

share the last piece in half (:

OpenStudy (anonymous):

Exactly so is 3 km is 20 minutes what is 1.5?

OpenStudy (anonymous):

If 3km = 20 min 1.5km =x Solve for x.

OpenStudy (ashleyy_smith):

3

OpenStudy (anonymous):

Incorrect, If 1.5 is half of 3 and 3km is 20 minutes, How long is 1.5 km?

OpenStudy (anonymous):

Time and distance are dependent on each other, If distance changes, the time does too, if the time changes, the distance does as well.

OpenStudy (anonymous):

Time and distance are dependent on each other, If distance changes, the time does too, if the time changes, the distance does as well.

OpenStudy (ashleyy_smith):

hmmmmmmmmmmmm

OpenStudy (anonymous):

So 3km = 20 minutes, If we chop 3 in half we get 1.5, Since the distance changed, The time has to also by the same amount. So 3km is now in half becoming 1.5km, what is 20 divided in half?